A unit square has a side length of 1 unit. Find the length of the diagonal of a unit square with a side length of 1 inch. Write your answer as a square root and to the nearest hundredth.
Answer by ewatrrr(24785)

(1)^2 + 1^2 = d^2
2 = d^2 (tossing out negative solution for unit measure)
√2 = d = 1.41 (to the nearest hundredth)